### v2

# The LCLStream Ecosystem for Multi-Institutional Dataset Exploration

Abstract.
We describe a new end-to-end experimental data streaming framework designed from the ground up to support new types of applications – AI training, extremely high-rate X-ray time-of-flight analysis, crystal structure determination with distributed processing, and custom data science applications and visualizers yet to be created. Throughout, we use design choices merging cloud microservices with traditional HPC batch execution models for security and flexibility. This project makes a unique contribution to the DOE Integrated Research Infrastructure (IRI) landscape. By creating a flexible, API-driven data request service, we address a significant need for high-speed data streaming sources for the X-ray science data analysis community. With the combination of data request API, mutual authentication web security framework, job queue system, high-rate data buffer, and complementary nature to facility infrastructure, the LCLStreamer framework has prototyped and implemented several new paradigms critical for future generation experiments.
Streaming, X-Ray Detectors, Data Analysis, Integrated Research Infrastructure
journalyear: 2026copyright: ccconference: SCA/HPCAsia 2026 Workshops: Supercomputing Asia and International Conference on High Performance Computing in Asia Pacific Region Workshops; January 26–29, 2026; Osaka, Japanbooktitle: SCA/HPCAsia 2026 Workshops: Supercomputing Asia and International Conference on High Performance Computing in Asia Pacific Region Workshops (SCA/HPCAsiaWS 2026), January 26–29, 2026, Osaka, Japandoi: 10.1145/3784828.3786261isbn: 979-8-4007-2328-5/2026/01ccs: Information systems Personalizationccs: Information systems Data encoding and canonicalizationccs: Information systems RESTful web servicesccs: Information systems Extraction, transformation and loading
1. Introduction
Autonomous experiment steering is already needed now to quickly adapt to changing experimental conditions that guide the instrument to optimal operating regimes. For example, machine-driven, Bayesian optimization of mechanical alignment settings for the incoming X-ray waveguides achieves 5x faster time-to-calibration than manual, human-driven optimization.(Roussel and others, 2024)
However, coupling High-Performance Computing (HPC) to external, on-line data sources requires the convergence of several new capabilities: data ontologies(Rajamohan and others, 2025) for the naming, processing, and storage of results at each level (e.g. raw events, X-ray crystal images, detected scattering peak positions, or electron detection times), experimental analysis frameworks for interactively guiding (CPU/GPU-intensive) data exploration, cross-facility coordination on co-scheduling of experiments and analysis, web-accessible applications programming interfaces (API-s) for data producers and HPC jobs, robust software for high-bandwidth data transmission, and a programming language coordinating interlinked, multi-participant activities.
This work discusses the design of a set of integrated services for experimental data collection and analysis that we developed. Together, they accomplish several objectives that would be impossible with a single, monolithic program: automated data collection, local fast data reduction, data streaming to academic or HPC facilities, user-defined analysis routines, and long-term data storage for data re-use (e.g. replicating studies or AI model training).
1.1. Data Streaming Framework
Experimental data at the Linac Coherent Light Sources (LCLS) is accessed using the psana1 and psana2 libraries (for LCLS and LCLS-II respectively).(Damiani and others, 2016; Thayer and others, 2024) These libraries read and assemble events from a running or archived experiment. These libraries are flexible, performant, and scalable across nodes using MPI, but must be used on site for administrative and technical reasons. The services that provide data access run user code and access databases (containing information needed to pre-process events). These actions can only be performed by authenticated users from within the LCLS facility.
The streaming framework described in this work opens up these data by providing local and remote distributed access to the psana1 and psana2 data outputs. Preliminary timing results show data arrival at an HPC job running on Oak Ridge’s National Center for Computational Sciences just seconds after collection at SLAC’s detectors in Menlo Park. The net result is increased data availability for the growing community of experts whose work focuses on experimental steering, time-sensitive processing, and online analysis.(Rakitin and others, 2020; Brim et al., 2024; Mariani and others, 2016)
Figure 1. Data streaming process diagram. Blue arrows show control paths, and black arrows show data flow. Dotted paths are for returned results. Event assembly and data formatting is performed by the psana framework on a compute job inside S3DF (left). The LCLStream API can start network buffers (on a S3DF data transfer node) and compute jobs on S3DF to format and send experimental data. External users can pair an LCLStream API call with jobs on other HPC clusters (right).
Fig. 1 shows the core components of our data streaming framework. The LCLStream API Server (in blue at the center of figure 1) allows beamline or external users to request a dataset from a specific experiment using a REST API with Javascript object notation (JSON)-formatted queries. The user performing the request and the server recognize each other through a certificate-driven mutual authentication process. LCLStream-API creates a JobID and launches a pair of processes on S3DF (the SLAC Shared Science Data Facility). First, it starts an NNG-Stream process a data transfer node with a new pair of recieve/send ports for this stream. Second, it starts LCLStreamer as an MPI job over SLURM on the S3DF cluster (on the left in the same figure). The receive URI is returned to the client. All compute processes can make independent connections to that address.
Together, the flexible data request API and the highly configurable LCLStreamer application give users flexibility in selecting datasets, performing partial data reduction, and choosing their preferred protocols and encodings. To support the data transfer, we have also created NNG-Stream (red box). It buffers data between parallel producers and consumers, smoothing the data flow in case of bursts of network capacity or activity. The buffer is stackable (not shown), so it can traverse complex network topologies. Data can be sent to multiple types of external applications – supercomputer centers, network-based appliances, and monitoring and automated control systems (in yellow on the right).
In the sections that follow, we give details on several experiments that were run using the early iterations of this infrastructure – including electron time-of-flight correlation analysis, image AI/ML training, and crystal structure inference from X-ray scattering. Then we go into greater detail on the interfaces, cache design, integration with job scheduling, and key authentication and performance characteristics. Overall, rapid progress has been made by keeping these individual components simple – speeding up the development cycle and decreasing the barrier for community involvement.
2. Scientific Applications
Application scientists are most interested in their specific problem domain. However, almost every beamline has a different way to access its output, along with metadata like detector positions, timestamps, and other instrument settings. Users often have little inclination to learn a new data format or data processing library. For these reasons, users almost always start their work by converting from the original data into their own formats. Although the original structure and ways of working with the data are lost, users are now in possession of an initial reduced dataset that is much more useful for their work.
Given this historical experience, the OM (OnDA Monitor) package for crystallographic data analysis has built a specification for a data pre-processing pipeline.(Mariani and others, 2016) In order to adapt this idea to LCLStreamer, we considered the needs of several new and existing data analysis projects:
  • MAXIE - a masked autoencoder for X-ray Image reconstruction,
  • PeakNet - an AI/ML method for peak-detection in crystal scattering X-ray images,
  • TMO-prefex - a data reduction and fast histogram/correlation counting package used during the commissioning of the LCLS-II upgrade for the Time-resolved atomic, Molecular and Optical Science (TMO) beamline.
  • CrystFEL - a program package for processing data collected in Serial Femtosecond Crystallography (SFX) experiments
Specifically, we looked for patterns in the data reduction steps and the input data formats used by each application. We also considered the needs of several developing projects which could make use of this data processing pipeline, including: online calibration estimation methods, high-rate image transmission from the X-Ray pump-probe (XPP) hutch to GPU nodes for image processing, and custom setups for CCTBX users who want more control over their data formatting (e.g. transformation to NeXus data format).
Each of these projects has unique output types: foundational AI model weights, reduced peak positions, histograms and experimental configuration-sensitive summary statistics, and electron density maps. However, all of them still bring event output and metadata through a data processing pipeline with well-defined processing steps (predictable compute requirements and intermediate result types).
At the beginning of our work, all of the above applications were able to run locally at the S3DF, by reading xtc or xtc2-formatted event data using LCLS’s psana package.(Damiani and others, 2016; Thayer and others, 2024) These event data files are output by each experimental run, can be generated at rates up to gigabytes per second (which will become terabytes in the future), and are effectively stored/loaded/streamed in parallel by LCLS’s high-throughput hardware setup. As the project progressed, we merged the common steps in these applications, and abstracted the differences into configurable options. Most variations can now be handled by adding new input detectors and data reduction functions, rather than rebuilding the entire data processing application.
2.1. Masked X-ray Image Autoencoder (MAXIE) and PeakNet
Cong and Chen describe two AI models capable of reconstructing and interpreting X-ray images.(Wang et al., 2025; Chen et al., 2025) The team has developed the Masked Autoencoder for X-ray Image Encoding (MAXIE), supporting model architectures ranging from hundreds of millions to billions of parameters. MAXIE is trained on approximately 286 terabytes of X-ray diffraction images from a variety of different detectors. Training epochs require hours to days depending on available compute resources, motivating its checkpointing and fault tolerance features. The implementation supports multiple parallelization strategies within a unified training framework, including single GPU, multi-GPU, and multi-node configurations using both Distributed Data Parallel (DDP) and Fully Sharded Data Parallel (FSDP) approaches (including sharded and full checkpoints), with optimizations including shared memory utilization and job scheduler integration for fault-tolerant execution.
The goal for this application is to run unsupervised training of this model on the full 286-terabyte dataset, potentially unlocking representations learned from the complete diversity of crystallographic data. Current production results come from a complementary supervised learning approach operating on a significantly smaller, algorithmically labeled dataset. Because both approaches leverage the same unified training infrastructure, scaling up to larger datasets is a matter of compute availability and network data accessibility. Architectural consistency in the AI/ML framework provides operational benefits including shared fault-tolerance mechanisms, unified checkpoint management, and the ability to scale supervised training to larger datasets as they become available.
2.2. TMO-prefex
The LCLS-II upgrade for the Time-resolved atomic, Molecular and Optical Science (TMO) beamline is designed to operate with X-ray laser shot repetition rates up to 1 MHz, collecting electron time of flight data with sub-femtosecond resolution after each shot.(Walter and others, 2021)
In theory, the increase in shot repetition rate from the previous 120 Hz allows experiments to complete 8333 times faster. However, the current state of the art for data analysis is running offline, out of the main experimental path. When done this way, it is not possible to determine whether each experimental step has been correctly configured and informative data has been gathered fast enough to keep up with the speed of data generation. In order to actually accelerate the experimental data collection phase, we need a fast data analysis pipeline. Speeding up the iteration time needed for completing each experimental step is key to realizing 1000x productivity increases. It will ultimately require data analysis capable of reconfiguring the experimental parameters in an automated way.(Roussel and others, 2024)
One central output of the TMO beamline is electron time of flight detection (FEX detector). Electrons emitted by molecules trapped in the TMO chamber are emitted at specific times after an initial laser excitation. Because one molecule can emit several electrons, the times and angles form a correlated signal, reporting on the molecule’s relaxation process.
Figure 2 shows the dataflow for signal acquisition and initial reduction from the electron time-of-flight detectors.(Gouin-Ferland et al., 2022) Downstream processing is then used to take these raw arrival times and spectra and accumulate histograms of the electron arrival times and light spectra, as well as perform more detailed sorting and chain-of-event reconstruction. These output histograms are the basis for Angle-resolved photo- and Auger–Meitner electron spectroscopy analyses (ARPES and ARAES, respectively).
Figure 2. TMO time of flight (ToF) detector configuration for detecting time and angular distribution of emitted electrons (reproduced from Ref. (Gouin-Ferland et al., 2022)). ToF spectrometer signals are processed by analog electronics before being digitized. After event detection, the central FPGA (circled with a red dotted line), forwards event features from all 8 peripheral FPGA-s on to the S3DF data processing pipeline. Instead of a custom code juggling event extraction with LCLS-II’s internal data path, LCLStream simplifies data extraction from this source down to a list of detector names and compression options.
The data stream output by the detectors is a current read-out for all angular channels at all times with femtosecond resolution. This extremely large data stream is eventually compressed into a list of individual electron arrival times. Three important intermediate steps are 1) raw waveform data, 2) a compressed set of waveforms above a threshold value, and 3) the final arrival times and detector numbers of current peaks.
Analysis on each type of data has historically been accomplished using ipython notebooks that parse data from hdf5-formatted dictionaries of arrays (several named arrays per detector). Even with good data tracking methods, it quickly becomes difficult to manage the combination of detector configurations, configuration options for processing steps, and conclusions made from analysis output. More importantly, this process has a built-in latency because it cannot be automated.
We incorporated the full data collection, transformation, and output to HDF5 pipeline into LCLStreamer by modeling the FEX detector, the three types of compression as data processing steps, and the output to HDF5 using an HDF5 data serializer. In our tests and evaluation of this framework, we simultaneously started 128 MPI-parallel data producers across 2 nodes of the S3DF data analysis cluster, an instance of NNG-Stream on a data transfer node of S3DF, and a 8 MPI-parallel data receivers across two nodes of OLCF’s ACE testbed.
The combination of steps above demonstrated important advantages in automation, metadata tracking, and workflow re-usability. Experimental configurations are documented in the run log curated by SLAC’s Elog system, as usual. Data processing pipeline is captured both by LCLStreamer’s configuration file and, when source changes are needed, by its git version control. Parallel processing, setup and tuned by facility staff, is used both for data output at S3DF and analysis at the OLCF HPC site. Finally, API-driven workflows substantially increase the accessibility and transparency of the above steps.
2.3. CrystFEL
CrystFEL (White et al., 2012) is a collection of programs for processing data collected using the Serial Femtosecond Crystallography technique. It includes programs performing all the steps needed to turn a set of diffraction images in a list of structure factors. This includes utilities to index and integrate diffraction patterns, and to merge measured intensities.
Live data analysis as the user is taking data at a beamline is a killer feature because it allows the user to maximize the scientific output from their beamtime.(Brewster et al., 2025) For example, determining whether detectors are aligned, diffraction spot intensities cover a large enough dynamic range, or whether the crystal itself contains growth defects strongly impact the timeline of an experiment.
In support of this, CrystFEL can operate both on images stored in HDF5 format(Group, 2025) in local files, or streamed via a network socket in the same format. Typically a data packet includes a diffraction image, and supplemental information needed for its interpretation (an estimation of the distance between the detector and the interaction point of the experiment, approximate beam energy, optical laser status, etc). Some additional pertinent experimental information (detector geometry, space group symmetry, etc.) is instead provided via program configuration files or command-line parameters. A graphical front-end to control the processing is available.
Previous works showing live, HPC-assisted data analysis mirrored raw (xtc) data files to the remote facility and then ran the full psana-based software load process.(Blaschke and others, 2024) In the present work, we aim to do the extraction and compression in parallel at the source. Separating the data pipeline this way is somewhat simplified by the program’s ability to read source data from the network. However, it is still challenging to arrange sending of the extra metadata while starting and stopping the program on an HPC center. If successful, the receiving analysis software will be less complicated and the entire pipeline will run with lower latency and higher throughput.
3. Software Infrastructure
In order to support the types of applications described above, we have constructed the LCLStream data pipeline from a modular set of network services.
  • LCLStreamer: An engine for fast, flexible data reduction and formatting
  • LCLStream-API: An HTTPS-REST API wrapping LCLStreamer and NNG-Stream
  • NNG-Stream: A buffer for receive-once, send-once message distribution from a concurrent set of producers to a concurrent set of consumers
  • Elog: LCLS’s electronic logbook, which enables run tracking and automated tasks
  • Psi-k and Psik-API: A portable batch submission interface for jobs with HTTPS-REST API job and file access
  • certified: An x.509 public key infrastructure for secure, mutually authenticated HTTPS client/server communication
LCLStreamer was designed based on patterns proven in the OM X-ray analysis software suite.(Mariani and others, 2016) This package has been developed over several years and has a strong user base within the X-ray science community. Its interface is designed around a top-level specification file for the data collection, reduction, and analysis pipeline. Users are able to achieve flexibility by plugging in different producers, data processing steps, and consumers within that pipeline.
The LCLStream API is a simple API wrapper starting and stopping LCLStream data producers. This enables off-site users to send an API request and then begin receiving data.
NNG-Stream is a message buffer using the nanomsg next generation library. It is designed to be simple and high-performance, operating by storing messages in memory and sending them in first-in-first-out order. It does no inspection of message contents.
Elog has been LCLS’s working database for many years. It tracks experiment metadata like run parameters, start and stop times, comments from beamline users, and other diagnostic and analysis steps performed as part of a given experiment. The usual workflow for experimental investigations involves setting up instrument parameters, starting and stopping data collection (creating a ”run”), and performing post-experiment annotation and analysis to the resulting run entry in the Elog. This operation makes Elog a natural place to store run-associated data, as it becomes a central location for organizing and automating runs forming an experiment and their associated run events.
Psi-k is a front-end to a file folder structure storing one job per folder. It creates new job folders from a JobSpec data structure (parsed from json or yaml). Job scripts can be created for different backends, including SLURM. Each job script runs psik reached to record its progress through a state sequence (queued $\to$ active $\to$ completed/canceled/failed).
The state sequence and the format of the data structures used by Psi-k benefit strongly from specifications developed by the ExaWorks PSI/J project.(Hategan-Marandiuc and others, 2023) Psi-k’s design diverged, however, in order to provide web interoperability. For example, Psi-k does not run job status checks by polling SLURM commands. State changes are stored in a status file, and can also trigger webhooks. Stdout and stderr are captured in a logs folder, and numbered sequentially for each re-run of the job. These file layouts form the basis for exposing jobs via an API (Psik-API). Mapping job create-read-update-destroy operations to the file hierarchy provides a straightforward REST-HTTPS interface to Psi-k.
3.1. Fast Data Reduction with LCLStreamer
LCLStreamer exports LCLS’s psana-native events (xtc/xtc2 format) to custom formats as needed by SLAC beamline users. It does this using the full parallelism supplied by LCLS’s data processing pipeline, applying all LCLS-related corrections and calibration. Key for our users, it has a flexible processing and formatting step. It outputs data suitable for direct consumption by the user’s preferred external software. Data from several events is accumulated, serialized, and finally passed to handlers to save as a file, network stream (or both) to external applications.
The LCLStreamer application is made up of several parts:
  • An Event Source (that generates the data)
  • A Processing Pipeline (that performs data reduction, calibration, etc.)
  • A Serializer (that turns the data into a binary object)
  • One or more Data Handlers (that write the binary object to a file, sends it through a network sockets, etc)
The LCLStreamer pipeline starts by retrieving a single event from an EventSource, and extracts from the event all data requested by the user. Any remaining data in the event is discarded. The data retrieved for each event has the format of a Python dictionary of Numpy Arrays. Each key in the dictionary corresponds to a data source. The value associated with the key is the information retrieved from the data source for the event being processed.
The operations of a ProcessingPipeline are then applied to the data retrieved from each event. This step makes use of the stream.py library to compose together a series of python generators (coroutines).
The standard pipeline batches together the results of processing several consecutive events. This accomplishes the same kind of batching one sees in a pytorch DataLoader, a key enabler of performance and interoperability. At that point, the accumulated data is returned in bulk. The data still has the format of a dictionary of Numpy array, with each key representing a data entry, and the corresponding value storing the accumulated data. The data is then serialized into a binary form.
Finally, the data is passed to one or more DataHandlers that can forward the data to the filesystem or any other external application, often via a network on in-process kernel socket, optionally in a compressed form. If multiple DataHandlers are present, they handle the same binary blob in parallel. New Serializers and DataHandlers can be added so that external applications can consume the data format they prefer.
Since there are multiple choices for event sources, processing pipelines, data serializers, etc., each section of the configuration file is marked with a type to select which implementation class it will use.
Next to the type selection, additional configuration options can also be provided to each of these implementation classes. For example, the HDF5Serializer class implements the data serializer interface. It serializes its input data into a binary string with the internal structure of an HDF5 file. Its configuration options can defined be when the data_serializer section selects the HDF5Serializer type.
data_serializer:
    type: HDF5Serializer
    compression_level: 3
    compression: zfp
    fields:
        timestamp: /data/timestamp
        detector_data: /data/data
Finally, the data_sources section of the configuration file defines the data that LCLStreamer extracts from every data event it processes. If a piece of information is part of the data event, but not included in the data_sources section, LCLStreamer will ignore it (filtering at read time). The data_sources section of the configuration file consists of a dictionary of data sources. Each entry’s key acts as a variable name that identifies the extracted data throughout the whole LCLStreamer data workflow. The entry’s associated dictionary defines the nature of the data source (via the mandatory type entry) and any other parameters needed to configure the initial data extraction from the psana library. The type of a data source is the name of a Python class within LCLStreamer that implements extraction.
For example, the following blocks declares variables timestamp and detector_data.
data_sources:
    timestamp:
        type: Psana1Timestamp
    detector_data:
        type: Psana1AreaDetector
        psana_name: Jungfrau1M
        calibration: true
The timestamp data class is of type Psana1Timestamp. Inside LCLStreamer, the Psana1Timestamp class will be called to read the associated timestamp data for each event. The detector_data class is instead of type Psana1AreaDetector. The two configuration parameters psana_name and calibration are passed to the Python class Psana1AreaDetector that defines how this type of data is retrieved.
LCLStreamer currently supports psana (Damiani and others, 2016) and psana2 (Thayer and others, 2024) as data sources, with a wide array of detectors and instrument readouts. It allows data to be serialized in customizable formats like HDF5 and can write them to files and/or send them via ZMQ or NNG sockets.
3.2. Data availability with LCLStream-API
LCLStream-API provides a web interface for external users to request data from LCLStreamer. It is built around LCLStreamer’s configuration file. As a REST-API, data transfers are started by POST operation, sending the configuration file as a typed JSON message to the transfers path. The response is either a validation error, or the ID for the newly created transfer. Issuing a GET or a DELETE to transfers/ID then reads the transfer status or stops a running transfer.
Although the API itself is simple, there are several implementation details that were very important to get right, including user authentication, parallel execution of LCLStreamer, and concurrent management of the message buffer. The certified package described below was used for user authentication, and the NNG-Stream package was used as a message buffer. A finite state machine was designed to ensure correctness of handling all the actions involved in the transfer process. State transitions for each transfer are driven by callbacks from the locally running NNG-Stream and the remotely running LCLStreamer, as well as user API calls to LCLStream-API. Section 3.4 describes the callback mechanism.
3.3. Message Buffering with NNG-Stream
To support LCLStreamer’s parallel data producers and consumers working, we need a message buffer capable of aggregating traffic from both sides at high speed. We built NNG-Stream based on the description of the Greta/Deletria forward buffer,(Cromaz et al., 2021) as their software was not available.
Placing a message buffer in-between these two provides several advantages over direct communication: no requirement for manually assigning producer/consumer address pairs, resilience to delays or crashes of individual producers or consumers, network traffic aggregation, and reduced network reachability requirements.
Figure 3. NNG-Stream Connectivity diagram. Each cache stores messages from all producers in a circular buffer, and distributes them round-robin to all consumers in an at-most-once fashion. Connectivity is provided via NNG Push0/Pull0 socket types and, optionally ZeroMQ Push sockets.(D’Amore, 2025) Multiple caches can work simultaneously to deliver traffic at rates on the order of tens of gigabytes per second.
Figure 3 shows data flow through NNG-Stream. Throughput tests run with a single cache on a laptop show aggregate bandwidth of 3 Gigabytes per second. These are limited only by local message routing and copying times. Network link speeds are quickly approaching the rate of 10s of Gigabytes per second. Hence, NNG-Stream, if replicated to 3 or 4 simultaneous caches, is capable of saturating these network links.
In our setup, NNG-Stream is run on a data transfer node at S3DF. These nodes serve incoming connections from both compute nodes inside S3DF as well as compute nodes at HPC facilities such as OLCF and NERSC. Direct outbound traffic is technically possible for most compute facility setups, but is sometimes blocked administratively. Nevertheless, this traffic path must to be supported to create a useful HPC facility working environment. Working around administrative blocks like this is possible by doing things like running another NNG-Stream on a data transfer node or data streaming node at the HPC facility. However, setting up these extra hops adds complexity and cost while providing little added security benefit.
Failure properties of the data stream are inherited from the underlying Push/Pull socket model. Producers and consumers can connect and disconnect from the cache without impacting the streaming status. Hence, producer or consumer process failures will result in dropping the data residing in that process. However, MPI usually terminates when a node fails, which can cause a cascade failure. A cache failure will terminate the stream, requiring a restart.
Normal stream shutdown is triggered by sender disconnect events. When all senders have disconnected, the cache enters a drain state, where no new producer connections are allowed. When all its data has been sent, the cache disconnects and exits. Clients are setup to detect this disconnect as an end-of-stream event. Alternatively, NNG-Stream can be configured to send ZeroMQ push data, and to send empty frames as sentinal values on stream end.
3.4. Experiment Management with E-Log, ARP and Airflow
At LCLS, the the Electronic Logbook (Elog) is the main entry point for users to track their experiments. It provides various services, such as recording when and how data were collected, logging user comments and attachments and managing a file catalog showing which files have been written. It also includes an Automated Run Processor (ARP) that allows to start data processing workflows without user interaction. In the Elog, users can define processing pipelines that are launched on specific events during the experiment (for example, when a data collection run begins or ends, when all files with the collected data are written to a particular storage location, etc.). The ability of the ARP to react to various different events is made possible by the components of the LCLS data management infrastructure exchanging information about their status via Apache Kafka messages.
The ARP is flexible, and can start different types of workflows tied to run events. The two most common cases are simple shell scripts that submit jobs to a batch queue, or sophisitcated worflows managed by an Apache Airflow instance. Special Airflow operators have been written to interact with the S3DF cluster via the SLURM task management system, and with the Perlmutter supercomputer at NERSC via the Superfacility API.(Enders and others, 2020) An additional operator that will allow communication with Psi-K (see 3.4) is currently under development, with the goal to allow Airflow to schedule tasks on a higher number of HPC facilities using different task scheduling systems.
Since the E-Log/ARP integration is operational, it is simple to enable an Airflow operator to manage an LCLStreamer-API transfer on experiment start/stop. This will start the data streaming pipeline as soon as a data collection run is started. Then compute-intensive analysis jobs will run remotely while the experiment is taking place.
3.5. HPC Job Management with Psi-K
The Psik and Psik-API projects provide a web-enabled interface for batch queuing systems on local machines and HPC centers. As described above, Psi-k is organized around a file layout where a top-level directory stores a collection of jobs. Documents within a job are rigidly structured. Files contained within each job’s folder (jobs/JobID) include: the JobSpec, a record of job state changes, job logs, and working directory files (as a user-managed collection).
As an example, the following Psi-k JobSpec shows how LCLStreamer can be deployed on S3DF’s data analysis cluster.
JobSpec:
    name: "lclstreamer"
    directory: "/psik/76312231.123/work"
    script:
      "mpirun -n120 lclstreamer -c cfg.yaml"
    resources:
      duration: 60 # minutes
      node_count: 1
      processes_per_node: 120
      cpu_cores_per_process: 1
    backend: S3DFslurm
    # POST to this URL on state change
    callback:
      "https://sdfdtn...edu/callbacks"
    cb_secret: "***"
This structure combines ideas from the ExaWorks PSI/J project(Hategan-Marandiuc and others, 2023) with traditional UNIX and cloud-based queuing systems. Note that PSI/J is primarily a script templating and shell-based polling engine, not designed around an asynchronous, web-API model. Hence, it was not suitable for this work.
Jobs are queued by a POST operation, sending a JobSpec json to the jobs path. The server responds with either a validation error or a new JobID. Issuing a GET or a DELETE to jobs/JobID then gets information about the job or cancels a queued or running job. The result is is a single-document method for launching HPC jobs.
API security is extremely important to design from the start. With this in mind, all communication with the API is strictly typed using data models, following the pydantic paradigm. In addition, more sensitive options like backend specification, wrapper scripts, and queue options are part of Psik-API’s offline configuration.
Internally, the server names and configures each backend (local, SLURM, API-client, etc.) using a BackendConfig, as shown below.
S3DFslurm: # BackendConfig
    type: slurm
    queue_name: milano
    project_name: lcls:tmox42619
Since each JobSpec names its backend, this extra information does not need to be a part of the API path (differing from NERSC’s Superfacility API(Enders and others, 2020)).
This simplifies both client and server configurations. API clients see all jobs together in a flat list, rather than separate for each backend. On the server side, backends are logical rather than physical. They may refer to different machines, partitions, or job scheduler attributes within a partition.
As a job, LCLStreamer is interesting because its primary activity is to send data over a network. This requires constant communication about the state of the activity. As Psi-k jobs proceed through execution phases, they update their state files and (optionally) send callbacks.
Here is an example callback sent to inform the LCLStream-API when the job above completes.
Callback:
    jobid: 76312231.123
    jobndx: 1
    state: completed
    info: ""
Similar callbacks are sent on cancellation or failure. These are used to manage the corresponding network data buffer run on the data transfer node.
In addition to the above network callbacks, the job also sends its stdout and stderr to logfiles. These can be read from the filesystem or, if the job is managed by Psik-API, via fetching or tailing the logfile.
Similar to Globus Compute,(Liu and others, 2021; Ananthakrishnan and others, 2024) the API listener can be run by individual users on a compute cluster’s login node. Alternatively, on Oak Ridge’s Frontier system, we run it on a Kubernetes pod with network access to the HPC cluster’s job scheduler and filesystem.(Papadimitriou et al., 2020) Oak Ridge is also implementing S3M API.(Skluzacek et al., 2025) As a client application, Psi-k can send jobs to this API, as well as to the NERSC Superfacility API(Enders and others, 2020) and Globus compute.(Ananthakrishnan and others, 2024) Differing from Globus compute, communications pass directly from the user to the Psik-API, bypasing a cloud services.
As an HTTPS server, user authentication for Psik-API can be handled several different ways. The simplest and most secure is to enable mutual TLS using the certified package. This requires signing user certificates, however, and not all centers are set up to do this at present. It is also possible to use a reverse proxy to access Psik-API and pass user details in the HTTP header. OLCF Slate uses this method, since Slate provides its own HTTPS termination and authentication scheme for users based on center-issued RSA tokens.
3.6. Mutual authentication with Certified
The Certified python package secures end-to-end communications with strong authentication for every client-server and server-server interaction. It addresses the central challenge of secure key distribution by providing public keys. The command-line certified program can create and sign a chain of x.509 certificates using ed25519 public keys and signatures. It is designed so that every python virtual environment maintains its own separate authentication and signing key. This way, an end-user, an application acting on the users behalf, a microservice, and an HPC facility can all have different keys that uniquely identify who is talking to who.
Trust in certified is established via digital signatures. Each client is expected to obtain signatures and a list of microservices from each organization they interact with. The client stores those signatures and microservice nicknames inside its configuration directory. Then, when the client wants to issue a message to a particular HTTPS microservice, it looks up the microservice URL and correct signature from its configuration directory.
Certified’s command-line interface makes these activities simple. Its documentation describes creating and signing certificates, as well as managing an individual list of named, trusted microservices. It also contains a wrapper to launch FastAPI web-servers.
Certified comes with a message command-line interface mimicking cURL for sending messages to REST-APIs. Although its functionality could be duplicated with cURL, performing the server URL translation and key lookup is complicated and error-prone.
A companion signer package is designed to issue signed user certificates on a UNIX/Linux system. It can be deployed with essentially zero configuratoin by a typical HPC facility on any login node. Similar to MUNGE (the scheme for issuing user authentication credentials for launching jobs within the SLURM job scheduler) it provides assertions about the user’s login name. In detail, it takes a certificate or a certificate signing request from a user, reads only the user’s public key, and issues the user a certificate linking their public key to their UNIX login name on that system. The user’s login name is determined by asking the kernel for the peer’s SO_PEERCRED information. Differing from MUNGE, it issues signed public keys. There is no danger of exposing these keys, since they can only be used by the original user who posesses the corresponding private key. Certified and signer never send the private key off of the user’s device.
Both certified and signer have strong logging integration. Certified provides a log-formatter that outputs the path accessed and the client’s identity in JSON formatting. It can optionally be configured to pass those logs to a Loki server for viewing within Grafana. The signer package stores its signatures to a database. Signature entries in its database can be queried for revocation status.
The combination of features above make certified a simple, viable, drop-in authentication method for developing microservices in an Integrated Research Infrastructure ecosystem.
4. Results
This section provides details on the use of LCLStream to address each of the application areas above. For all of the areas, the main result to note was that the framework successfully re-formatted LCLS data for remote consumption and simultaneously transmitted the data over the network to a compute job running on OLCF. The round-trip network latency measured between S3DF’s data transfer nodes and OLCF was consistently around 33-36 milliseconds, only about 50% longer than it takes light to travel the same distance.
All setups made use of parallel producers and parallel consumers. In most cases the bottleneck (determining maximum data throughput) was the data read/formatting speed at S3DF, around 1-3 GB/sec. In other cases, the cache was the rate-limiting step, at around 3-4 GB/sec. Running parallel caches could address the latter, but was not tested in this work.
4.1. Masked X-ray Image Autoencoder (MAXIE) and PeakNet
LCLS operates a variety of different detector types. Thus, the most complicated step in this workflow is curating the input data for training. We produced a series of LCLStreamer input files targeting various experiment/run IDs containing epix10k2M images from the Psana1AssembledAreaDetector source. We also included the photon wavelength read from psana’s SIOC:SYS0:ML00:AO192 as annotation. We opted to implement a special-purpose “PeaknetPreprocessingPipeline” to center and pad images to consistent sizes on the data generation side.
In this application, the LCLStreamer input files replaced a custom data extraction code. The result ended up being more readable and helped with documentation and re-use of the curation steps. Two primary difficulties were noted. First, at this stage of development, it is necessary to manually start the API on S3DF’s data transfer node. Second, we needed to implement our own client-side caching mechanism to prevent re-downloading data. This is significant for this application, since ML training makes many passes over its input. Still, the ability to download needed inputs on-demand decreased the data-gathering lag time between getting an HPC allocation and being able to run scientific work. It also eliminated stale data on the HPC side as a potential source of errors. These results suggest that we need robust, flexible methods like LCLStreamer for pre-processing our data as we work to improve its quality.
4.2. TMO-prefex
The TMO-prefex application was an early version of LCLStreamer. It was used during the commissioning of the TMO end-station just after its upgrade to LCLS-II’s higher data rate. We used Psik-API to run a job computing electron angular and time-of-flight correlation plots on OLCF ACE. Then we manually started MPI-parallel TMO-prefex data producers pushing data through NNG-Stream on S3DF. We synchronized work by interacting through the control room’s live Zoom channel. Due to compression at the source (time-windowed signal thresholding), we did not have a bandwidth limitation. The major source of startup-latency was the program start-up time for parallel data producers on S3DF.
4.3. CrystFEL
CrystFEL has added the capability to work on live data streamed from DECTRIS-type detectors. The data format is specified as part of their Simplon-API, and consists of both control packets and data packets. We implemented this by adding the SimplonBinarySerializer type as an alternative to HDF5 serialization. This serializer inserts the appropriate control messages into the output stream.
During a beamtime at the Macromolecular Femtosecond Crystallography (MFX) endstation at LCLS, the LCLStreamer component was used to send data from one of the experimental runs over to the testbed HPC cluster at the Oak Ridge Leadership Computing Facility, where CrystFEL had been setup to receive and process the transferred data. The latency between data collection and processing with CrystFEL has been shown to be within the range of 15-25 seconds, allowing useful feedback information within a time range sufficient for researchers to manually steer the experiment in the desired direction.
This application especially demonstrated the benefits obtained from standards combined with a modular, interoperable software design. We did not need to create a bespoke data pipeline to generate this message format specifically, nor did CrystFEL have to add a facility-specific data ingestion framework. Instead, we implemented only the specific data format, named it after the standard, and re-used the rest of the facility and user community software pipelines.
5. Discussion
Data streaming is an emerging technological area within DOE’s IRI (Integrated Research Infrastructure). Our work follows the pattern of the GRETA/Deleria project,(Cromaz et al., 2021) but uses published, open-source, and modular software components targeted to the field of X-ray image analysis. Compared to the edge-to-exascale framework,(Yin et al., 2024) our work focuses on well-defined intermediate data formats and repeatable, modular processing steps. We also target GB/sec data rates arising from high-frequency X-ray sources, as opposed to somewhat lower data rates from synchrotrons or MB/sec data rates from neutron detectors and automated laboratories. The latter types of data flows can still be handled by pub/sub message transport models, as demonstrated by INTERSECT(Thakur et al., 2022) or BlueSky(Wijesinghe et al., 2024) projects. Both the latter frameworks have the somewhat larger goal of controlling online experiments. We do not seek to replace those frameworks, but rather to provide a modular data request and transport mechanism.
Central to this project’s success has been a continued, close collaboration between the DOE-IRI pathfinder projects, LCLS, S3DF, and leadership compute facilities. The design boundaries between the components we eventually arrived at were shaped by sharing ideas and preliminary implementations of workflows among all these groups. The team-developed microservices described here work in concert to accomplish streaming. Continued work on the IRI progam promises even greater integration with facility-deployed infrastructure that will continue to improve their re-usability across sites and reliability for production work.
Metrics for measuring the pace of improvement include: 1) eliminating manual intervention steps during install, startup, and execution (such as starting servers and creating ssh tunnels) 2) reducing site-specific adaptations needed to interoperate across different experiment and compute facility API-s, 3) standardizing security technologies (such as mutual TLS authentication) needed to connect network sockets between data producers and consumers, 4) solving the ”co-scheduling” problem for simultaneous commitment of beam time, network bandwidth, and compute resources, and 5) providing unencumbered access to fast, open-source data movement services (e.g. XRootD or FTS3) for managing site-to-site file copies.
6. Conclusion
We have shown a new end-to-end experimental data streaming framework incorporating key design choices for security and flexibility. User interactions with API-s are authenticated and encrypted, and make use of strongly typed data schemas.
It has been designed from the ground up to support new types of applications – AI training, extremely high-rate X-ray time-of-flight analysis, crystal structure determination with distributed processing, and custom data science applications and visualizers yet to be created. These are supported by a modular software stack, where individual applications run as communicating microservices.
This project has a unique position in relation to other efforts within the DOE Integrated Research Infrastructure (IRI) landscape.(Rogers, 2024) We have chosen to focus on small services that collaborate to provide a high-rate data channel. This avoids dealing with the full complexity of the experimental orchestration layer that Bluesky,(Wijesinghe et al., 2024) INTERSECT,(Brim et al., 2024) and GRETA/DELERIA’s(Cromaz et al., 2021) Janus framework do. At the same time, the certified, psik, and NNG-Stream packages have given general solutions to microservice communication, HPC job submission, and high-rate data buffering. They are targeted to simple, drop-in operations on HPC clusters, and avoid a centralized, cloud controller.
High-speed data streaming is a significant need for the X-ray science community. The LCLStreamer framework has prototyped and implemented several new paradigms that meet this need. Its successful deployment depends on the collaboration and ongoing dedication of experts in experimental, compute, cloud service, and networking infrastructure.
Acknowledgements.
We thank Daniel Pelfrey, Ross Miller, Jordan Webb, Jordan Brown, Carl Bai, and Matthew Ezell for configuring high-speed network access from the Defiant system on OLCF’s ACE testbed and from compute nodes on OLCF Frontier. This research used resources of the SLAC National Accelerator Laboratory and the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. Use of the Linac Coherent Light Source (LCLS), SLAC National Accelerator Laboratory, is supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under Contract No. DE-AC02-76SF00515.